# Localization phenomena in the random XXZ spin chain

### v1

@@ 3. Basic features of the XXZ spin chain
We will denote by   $d_{1}(\mathbf{x},\mathbf{y})$ the $\ell^{1}$ distance between ${\bf x}$ and ${\bf y}$ in $\mathbb{Z}^{N}$, that is,
$d_{1}(\mathbf{x},\mathbf{y})=\left\lvert{\bf x}-{\bf y}\right\rvert_{1}=\sum_{i=1}^{N}\left\lvert x_{i}-y_{i}\right\rvert.$ (3.7)
In view of (2.11), we also introduce a modified $\ell^{1}$ distance between subsets of $\mathbb{Z}$:
$\widetilde{d}_{1}(\mathbf{x},\mathbf{y})=\begin{cases}d_{1}(\mathbf{x},\mathbf{y})&\quad\text{if}\quad\left\lvert\mathbf{x}\right\rvert=\left\lvert\mathbf{y}\right\rvert\\ \infty&\quad\text{otherwise}\end{cases}.$ (3.8)
An important property of Green functions is the Combes-Thomas bound [EKS1, Proposition 4.1],
$\sup_{z\in\mathds{H}_{m}}\left\lvert G^{\Lambda}_{z}({\bf x},{\bf y})\right\rvert\leq C_{m}\mathrm{e}^{-c_{m}\widetilde{d}_{1}(\mathbf{x},\mathbf{y})}\quad\text{for}\quad m<\tfrac{3}{4}\quad\text{and}\quad\mathbf{x},\mathbf{y}\in\mathcal{P}_{+}(\Lambda).$ (3.9)
@@
[EKS1, Proposition 4.1] gives a Combes-Thomas bound for $H^{\Lambda}$ with constants independent of $N$ and $\Lambda$, so combined with the definition (3.8) it clearly implies (3.9).
As mentioned in Section 1, $H^{\Lambda}_{N}$ is a random Schrödinger operator on a certain subgraph of $\Lambda^{N}$, but the standard Combes-Thomas bounds will only yield (3.9) with $N$ dependent constants. Although [EKS1, Proposition 4.1] is stated and proven on each $\mathcal{H}_{\Lambda}^{\left(N\right)}$, the proof uses the special structure of $H^{\Lambda}_{N}$ to obtain constants independent of $N$.
Unfortunately, we cannot use the Combes-Thomas bound directly for energies lying in $\mathds{H}_{m}$ for $m\geq\frac{3}{4}$. The way around is to lift the spectrum of the operator $H^{\Lambda}$, and note that the Combes-Thomas bound holds for the lifted operator.
@@
and recall that [EK22, Lemma 3.5]
@@
$\widehat{H}_{q}^{\Lambda}=\begin{cases}H^{\Lambda}+\left(1-\tfrac{1}{\Delta}\right)Q_{0}^{\Lambda}&\quad\text{if}\quad q=0,\tfrac{1}{2}\\ H^{\Lambda}+{\lceil q\rceil}\left(1-\tfrac{1}{\Delta}\right)\widehat{Q}_{\leq\lceil q\rceil}^{\Lambda}&\quad\text{otherwise}\end{cases}.$ (3.13)
The basic feature of these operators is that we have
$\widehat{H}_{q}^{\Lambda}\geq\left(1-\tfrac{1}{\Delta}\right)\;\text{for}\;q=0,\tfrac{1}{2}\quad\text{and}\quad\widehat{H}_{q}^{\Lambda}\geq\left({\lceil q\rceil}+1\right)\left(1-\tfrac{1}{\Delta}\right)\quad\text{otherwise},$ (3.14)
$\widehat{H}_{q}^{\Lambda}-E\geq\tfrac{1}{4}\left(1-\tfrac{1}{\Delta}\right)\quad\text{for all}\quad E\in I_{\leq{q}}\;.$
@@
The Combes-Thomas bound of [EKS1, Proposition 4.1] then holds for the modified Green functions: For all $q\in\frac{1}{2}\mathbb{N}^{0}$ we have
$\sup_{z\in\mathds{H}_{q}}\left\lvert\widehat{G}^{\Lambda}_{q,z}({\bf x},{\bf y})\right\rvert\leq C_{q}\mathrm{e}^{-c_{q}\widetilde{d}_{1}(\mathbf{x},\mathbf{y})}\quad\text{for all}\quad{\mathbf{x},\mathbf{y}\in\mathcal{P}_{+}(\Lambda).}$ (3.16)
@@
$\pi^{S}_{\nu}=\lceil q\rceil\left(1-\tfrac{1}{\Delta}\right)\widehat{R}^{S}_{q,\nu}\widehat{Q}^{S}_{\leq\lceil q\rceil}\pi^{S}_{\nu}=\lceil q\rceil^{2}\left(1-\tfrac{1}{\Delta}\right)^{2}\widehat{R}^{S}_{q,\nu}\widehat{Q}^{S}_{\leq\lceil q\rceil}\pi^{S}_{\nu}\widehat{R}^{S}_{pq,\nu}\widehat{Q}^{S}_{\leq\lceil q\rceil}.$ (3.17)
Since our arguments rely on a certain decoupling idea, we need to introduce yet another variety of Green functions. Given $K\subset\Lambda$, we consider the operator $H^{K}=H^{K}\otimes\mathds{1}_{\mathcal{H}_{K^{c}}}$ acting on $\mathcal{H}_{\Lambda}$. Then the decoupled Hamiltonian and resolvent on $\mathcal{H}_{\Lambda}$ are given by
$H^{K,K^{c}}=H^{K}+H^{K^{c}},\quad R_{z}^{K,K^{c}}=\left(H^{K,K^{c}}-z\right)^{-1},\quad\Gamma^{K}=H^{\Lambda}-H^{K,K^{c}}.$ (3.18)
The corresponding decoupled Green function is then
$G^{K,K^{c}}_{z}({\bf x},{\bf y})=\langle\phi_{\bf x},R_{z}^{K,K^{c}}\phi_{\bf y}\rangle.$ (3.19)
The corresponding modified Hamiltonian is $\widehat{H}^{K,K^{c}}_{m}=\widehat{H}^{K}_{m}+\widehat{H}^{K^{c}}_{m}$ acting on $\mathcal{H}_{\Lambda}$, and the modified Green function,
$\widehat{G}^{K,K^{c}}_{m,z}({\bf x},{\bf y})=\langle\phi_{\bf x},\widehat{R}_{m,z}^{K,K^{c}}\phi_{\bf y}\rangle,$ (3.20)
satisfies (3.16) as well. We note that $G^{K,K^{c}}_{z}({\bf x},{\bf y})$ (and its modified analogue) vanishes unless $\left\lvert{\bf x}\cap K\right\rvert=\left\lvert{\bf y}\cap K\right\rvert$ and $\left\lvert{\bf x}\cap K^{c}\right\rvert=\left\lvert{\bf y}\cap K^{c}\right\rvert$.
The proof of Theorem 2.4 will be facilitated by the following a-priori estimate (see, e.g., [EK22, Lemma 3.4]):
@@
where $\left\lVert\cdot\right\rVert_{2}$ denotes the Hilbert-Scmidt norm, which implies that
@@
We note that for real valued $z$, $G^{\Lambda}_{z}({\bf x},{\bf y})$ is understood here and below as $G^{\Lambda}_{z+i0}({\bf x},{\bf y})$.
If $\mathbf{x}\subset\Lambda$ and $S\subset\Lambda$, we write $\mathbf{x}_{S}=\mathbf{x}\cap S$. If $P$ is an orthogonal projection, we write $\bar{P}=\mathds{1}-P$.
@@
In this Section we prove Theorem 2.4. We fix $\Delta>1$, $\lambda>0$, and $s\in(0,\frac{1}{3})$.
We will use the following lemma. For $k\in\left\{1,2,\ldots,N\right\}$ we let
@@
Let $q\ \in\frac{1}{2}\mathbb{N}$, $1\leq q$, and $N\in\mathbb{N}$. Fix $\Lambda\subset\mathbb{Z}$, and and suppose (2.23) holds for all ${\bf x},{\bf y}\in\mathcal{P}_{N,\lceil q\rceil}(\Lambda)$. Then (2.23) holds for all ${\bf x},{\bf y}\in\mathcal{P}_{N}(\Lambda)$ (with different constants, independent of $\Lambda$ and $N$).
@@
$\sup_{z\in\mathds{H}_{q}}\sup_{\Lambda\subset\mathbb{Z}}\mathbb{E}_{\Lambda}\left\{\left\lvert G^{\Lambda}_{z}({\bf x},{\bf y})\right\rvert^{s}\right\}$ $\leq C_{q}\mathrm{e}^{-c_{q}\left\lvert{\bf x}-{\bf y}\right\rvert_{1}}+C_{q}\sum_{{\bf u}\in\mathcal{P}_{N,\lceil q\rceil}(\Lambda)}\mathrm{e}^{-c_{q}\left\lvert{\bf x}-{\bf u}\right\rvert_{1}}\mathrm{e}^{-c_{q}\left\lvert{\bf u}-{\bf y}\right\rvert_{1}}$ (4.4)
$\quad\quad+C_{q}\sum_{{\bf u},{\bf v}\in\mathcal{P}_{N,\lceil q\rceil}(\Lambda)}\mathrm{e}^{-c_{q}\left\lvert{\bf x}-{\bf u}\right\rvert_{1}}\mathrm{e}^{-c_{q}d^{\Lambda}_{H}\left({\bf u},{\bf v}\right)}\mathrm{e}^{-c_{q}\left\lvert{\bf v}-{\bf y}\right\rvert_{1}}$
@@
where in the last step we used properties of exponential sums, see (B.2) below. ∎
@@
We first prove the lemma for $\mathbf{x},\mathbf{y}\in\mathcal{P}_{N,\lceil q\rceil}(D)$. This is done using the reduction to resolvents achieved by using the estimate [AW, Eq. (7.44)] and the spectral averaging as in [AW2, Theorem 4.5]. The final result can be re-formulated in our setting as:
@@
[…]
Note that that $d^{D}_{H}({\bf u},{\bf y})\geq d^{D}_{H}({\bf x},{\bf y})$ if $x\in\mathbf{u}\subset D$.
Given $\mathbf{x},\mathbf{y}\in\mathcal{P}_{N,\lceil q\rceil}$, we estimate $\mathbb{E}_{D}\left\{\mathcal{Q}_{q}^{D}({\bf x},{\bf y})\right\}$ by (4.8), and estimate the term $\mathbb{E}_{D}\left\{\left\lvert G_{E}^{D}({\bf u},{\bf y})\right\rvert^{s}\right\}$ inside the integral as in (4.4), using (4.5) , getting
$\mathbb{E}_{D}\left\{\mathcal{Q}_{q}^{D}({\bf x},{\bf y})\right\}$ $\leq C_{q}\sum_{{\bf u}\in\mathcal{P}_{N}(D):x\in{\bf u}}\mathrm{e}^{-c_{q}\left\lvert{\bf u}-{\bf y}\right\rvert_{1}}+C_{q}\sum_{{\bf u}\in\mathcal{P}_{N}(D):x\in{\bf u}}\ \sum_{{\bf v}\in\mathcal{P}_{N,\lceil q\rceil}(D)}\mathrm{e}^{-c_{q}\left\lvert{\bf u}-{\bf v}\right\rvert_{1}}\mathrm{e}^{-c_{q}\left\lvert{\bf v}-{\bf y}\right\rvert_{1}}$ (4.9)
$\quad\quad+C_{q}\sum_{{\bf u}\in\mathcal{P}_{N}(D):x\in{\bf u}}\ \sum_{{\bf v},{\bf w}\in\mathcal{P}_{N,\lceil q\rceil}(D)}\mathrm{e}^{-c_{q}\left\lvert{\bf u}-{\bf v}\right\rvert_{1}}\mathrm{e}^{-c_{q}d^{D}_{H}\left({\bf v},{\bf w}\right)}\mathrm{e}^{-c_{q}\left\lvert{\bf w}-{\bf y}\right\rvert_{1}}.$
To bound the first sum, we note that
$\left\lvert{\bf u}-{\bf y}\right\rvert_{1}\geq d_{H}({{\bf u},{\bf y}})\geq\operatorname{dist}_{D}(x,{\bf y})=d_{H}^{D}({\bf x},{\bf y}),$ (4.10)
using (4.7). Hence
$\sum_{{\bf u}\in\mathcal{P}_{N}(D):x\in{\bf u}}\mathrm{e}^{-c_{q}\left\lvert{\bf u}-{\bf y}\right\rvert_{1}}\leq\mathrm{e}^{-\frac{c_{q}}{2}d_{H}^{D}({\bf x},{\bf y})}\sum_{{\bf u}\in\mathcal{P}_{N}(D)}\mathrm{e}^{-\frac{c_{q}}{2}\left\lvert{\bf u}-{\bf y}\right\rvert_{1}}\leq C_{q}\mathrm{e}^{-\frac{c_{q}}{2}d_{H}^{D}({\bf x},{\bf y})},$ (4.11)
@@
$\left\lvert{\bf u}-{\bf v}\right\rvert_{1}+\left\lvert{\bf v}-{\bf y}\right\rvert_{1}$ $\geq\left\lvert{\bf u}-{\bf y}\right\rvert_{1}\geq d_{H}^{D}({\bf x},{\bf y}),$ (4.12)
$\left\lvert{\bf u}-{\bf v}\right\rvert_{1}+\left\lvert{\bf v}-{\bf y}\right\rvert_{1}$ $\geq\tfrac{1}{2}\left(\left\lvert{\bf u}-{\bf y}\right\rvert_{1}+\left\lvert{\bf v}-{\bf y}\right\rvert_{1}\right).$
@@
$\sum_{{\bf u}\in\mathcal{P}_{N}(D):x\in{\bf u}}\ \sum_{{\bf v}\in\mathcal{P}_{N,\lceil q\rceil}(D)}\mathrm{e}^{-c_{q}\left\lvert{\bf u}-{\bf v}\right\rvert_{1}}\mathrm{e}^{-c_{q}\left\lvert{\bf v}-{\bf y}\right\rvert_{1}}$ (4.13)
[…]

### v3

@@ 3. Basic features of the XXZ spin chain
Given $N\in\mathbb{N}$, we identify $\mathbf{x}\in\mathcal{P}_{N}(\Lambda)$ with $(x_{1},\ldots,x_{N})\in\Lambda^{N}$, where $x_{1}<\ldots<x_{N}$. We introduce a distance in $\mathcal{P}_{N}(\Lambda)$ by
$d_{1}^{\Lambda}(\mathbf{x},\mathbf{y})=\sum_{i=1}^{N}\operatorname{dist}_{\Lambda}(x_{i},y_{i})\quad\text{for}\quad\mathbf{x},\mathbf{y}\in\mathcal{P}_{N}(\Lambda).$ (3.7)
We may have $d_{1}^{\Lambda}(\mathbf{x},\mathbf{y})=\infty$ if $\Lambda$ is disconnected. Note that
$d_{1}^{\Lambda}(\mathbf{x},\mathbf{y})$ $\geq d_{1}(\mathbf{x},\mathbf{y})=d_{1}^{\mathbb{Z}}(\mathbf{x},\mathbf{y})=\left\lvert{\bf x}-{\bf y}\right\rvert_{1}=\sum_{i=1}^{N}\left\lvert x_{i}-y_{i}\right\rvert$, (3.8)
$d_{1}^{\Lambda}(\mathbf{x},\mathbf{y})$ $\geq d_{H}^{\Lambda}(\mathbf{x},\mathbf{y}).$
In view of (2.11), we also introduce a distance on $\mathcal{P}_{f}(\Lambda)$ by
[…]
Since our arguments rely on a certain decoupling idea, we need to introduce graphs obtained from $\Lambda$ by decoupling. Given $K\subset\Lambda$ (we allow $K=\emptyset$), we consider the graph $\Lambda^{\left(K\right)}$ with vertex set $\Lambda$ and edges totally contained in either $K$ or $K^{c}=\Lambda\setminus K$, so $K$ is disconnected from $K^{c}$. (In particular, $d_{\Lambda^{\left(K\right)}}(K,K^{c})=\infty$.) We define $d_{1}^{K,K^{c}}(\mathbf{x},\mathbf{y})=d_{1}^{\Lambda^{\left(K\right)}}(\mathbf{x},\mathbf{y})$ and $\widetilde{d}_{1}^{K,K^{c}}(\mathbf{x},\mathbf{y})=\widetilde{d}_{1}^{\Lambda^{\left(K\right)}}(\mathbf{x},\mathbf{y})$ as in (3.7) and (3.9) using the distance $d_{\Lambda^{\left(K\right)}}(\cdot,\cdot)$. Note that $\widetilde{d}_{1}^{\Lambda^{\left(K\right)}}(\mathbf{x},\mathbf{y})\geq\widetilde{d}_{1}^{\Lambda}(\mathbf{x},\mathbf{y})$ for all $\mathbf{x},\mathbf{y}\in\mathcal{P}_{f}(\Lambda)$.
We consider the operators $H^{K}=H^{K}\otimes\mathds{1}_{\mathcal{H}_{K^{c}}}$ and $H^{K^{c}}=\mathds{1}_{\mathcal{H}_{K}}\otimes H^{K^{c}}$acting on $\mathcal{H}_{\Lambda}$. Then the decoupled Hamiltonian, resolvent, and boundary operator on $\mathcal{H}_{\Lambda}$ are given by
$H^{\Lambda^{\left(K\right)}}=H^{K,K^{c}}=H^{K}+H^{K^{c}},\;R_{z}^{\Lambda^{\left(K\right)}}=R_{z}^{K,K^{c}}=\left(H^{K,K^{c}}-z\right)^{-1},\;\Gamma^{K}=H^{\Lambda}-H^{K,K^{c}}.$ (3.10)
The corresponding decoupled Green function is then
$G^{\Lambda^{\left(K\right)}}_{z}({\bf x},{\bf y})=G^{K,K^{c}}_{z}({\bf x},{\bf y})=\langle\phi_{\bf x},R_{z}^{K,K^{c}}\phi_{\bf y}\rangle.$ (3.11)
Note that $G^{K,K^{c}}_{z}({\bf x},{\bf y})$ vanishes unless $\left\lvert{\bf x}\cap K\right\rvert=\left\lvert{\bf y}\cap K\right\rvert$ and $\left\lvert{\bf x}\cap K^{c}\right\rvert=\left\lvert{\bf y}\cap K^{c}\right\rvert$.
Since on $\operatorname{Ran}\chi_{\mathbb{N}}(\mathcal{N}^{\Lambda})$ we have
$H^{\Lambda^{\left(K\right)}}-E\geq\left(1-\tfrac{1}{\Delta}\right)\mathcal{W}^{\Lambda^{\left(K\right)}}-\tfrac{3}{4}\left(1-\tfrac{1}{\Delta}\right)\geq\tfrac{1}{4}\left(1-\tfrac{1}{\Delta}\right)\mathcal{W}^{\Lambda^{\left(K\right)}}\;\text{for}\;E\in I_{\leq\frac{1}{2}},$ (3.12)
we have, using the resolvent identity, that
$\sup_{z\in\mathds{H}_{\frac{1}{2}}}\left\lVert\left(\mathcal{W}^{\Lambda^{\left(K\right)}}\right)^{\frac{1}{2}}R_{z}^{\Lambda^{\left(K\right)}}\left(\mathcal{W}^{\Lambda^{\left(K\right)}}\right)^{\frac{1}{2}}\right\rVert\leq 2\sup_{E\in I_{\leq\frac{1}{2}}}\left\lVert\left(\mathcal{W}^{\Lambda^{\left(K\right)}}\right)^{\frac{1}{2}}R_{E}^{\Lambda^{\left(K\right)}}\left(\mathcal{W}^{\Lambda^{\left(K\right)}}\right)^{\frac{1}{2}}\right\rVert\leq\tfrac{8}{1-\frac{1}{\Delta}}<\infty.$ (3.13)
The estimate (3.13) is the required hypothesis for an important property of Green functions, a Combes-Thomas bound:
$\sup_{z\in\mathds{H}_{m}}\left\lvert G^{\Lambda^{\left(K\right)}}_{z}({\bf x},{\bf y})\right\rvert\leq C_{m}\mathrm{e}^{-c_{m}\widetilde{d}_{1}^{\Lambda^{\left(K\right)}}(\mathbf{x},\mathbf{y})}\quad\text{for}\quad m\leq\tfrac{1}{2}\quad\text{and}\quad\mathbf{x},\mathbf{y}\in\mathcal{P}_{+}(\Lambda),$ (3.14)
which holds for all $K\subset\Lambda$.
The bound (3.14) is proven similarly to [15, Proposition 4.1], which is stated for $H^{\Lambda^{\left(K\right)}}_{N}$ with $N\in\mathbb{N}$, and yields decay in the distance $d_{1}^{\mathbb{Z}}(\mathbf{x},\mathbf{y})=\left\lvert\mathbf{x}-\mathbf{y}\right\rvert_{1}$. To get the bound (3.14), in the proof we replace the distance $d_{1}^{\mathbb{Z}}(\mathbf{x},\mathbf{y})$ by $d_{1,R}^{\Lambda^{\left(K\right)}}(\mathbf{x},\mathbf{y})=\min\left\{d_{1}^{\Lambda^{\left(K\right)}}(\mathbf{x},\mathbf{y}),R\right\}$, where $R\gg 1$, perform the proof, and let $R\to\infty$. We get the bound (3.14) for $H^{\Lambda^{\left(K\right)}}_{N}$ with constants independent of $N$, $\Lambda$ and $K$, so combining with the definition (3.9) we get (3.14).
@@
[…]
Unfortunately, we cannot use the Combes-Thomas bound given in (3.14) directly for energies lying in $\mathds{H}_{m}$ for $m\geq 1$. The way around this obstacle is to lift the spectrum of the operator $H^{\Lambda^{\left(K\right)}}$, and show that (3.14) holds for the lifted operator.
To simplify notation, for the remainder of this section we will simply write $H^{\Lambda}$ for $H^{\Lambda^{\left(K\right)}}$, with the results holding for $H^{\Lambda^{\left(K\right)}}$ for all $K\subset\Lambda$.
@@
and recall that [12, Lemma 3.5]
@@
[…]
@@
These operators satisfy
$\widehat{H}_{q}^{\Lambda}\geq\left(1-\tfrac{1}{\Delta}\right)\;\text{for}\;q=0,\tfrac{1}{2},\quad\text{and}\quad\widehat{H}_{q}^{\Lambda}\geq\left({\lceil q\rceil}+1\right)\left(1-\tfrac{1}{\Delta}\right)\quad\text{otherwise},$ (3.20)
and, moreover, on $\operatorname{Ran}\chi_{\mathbb{N}}(\mathcal{N}^{\Lambda})$ we have
$\widehat{H}_{q}^{\Lambda}-E\geq\tfrac{1}{4\left({\lceil q\rceil}+1\right)}\left(1-\tfrac{1}{\Delta}\right)\mathcal{W}^{\Lambda}\quad\text{for all}\quad E\in I_{\leq{q}}\;\text{for}\;q\in\tfrac{1}{2}\mathbb{N}^{0},$ (3.21)
so it follows that
$\sup_{z\in\mathds{H}_{m}}\left\lVert\left(\mathcal{W}^{\Lambda}\right)^{\frac{1}{2}}\widehat{R}^{\Lambda}_{q,z}\left(\mathcal{W}^{\Lambda}\right)^{\frac{1}{2}}\right\rVert\leq 2\sup_{E\in I_{\leq\frac{1}{2}}}\left\lVert\left(\mathcal{W}^{\Lambda}\right)^{\frac{1}{2}}\widehat{R}^{\Lambda}_{q,E}\left(\mathcal{W}^{\Lambda}\right)^{\frac{1}{2}}\right\rVert\leq\tfrac{8\left({\lceil q\rceil}+1\right)}{1-\frac{1}{\Delta}}<\infty.$ (3.22)
The bound (3.21) is just (3.12) for $q=0,\frac{1}{2}$. To derive (3.21) for $q\geq 1$ on $\operatorname{Ran}\chi_{\mathbb{N}}(\mathcal{N}^{\Lambda})$, we use
$\widehat{H}_{q}^{\Lambda}-E$ $\geq\left(1-\tfrac{1}{\Delta}\right)\mathcal{W}^{\Lambda}+{\lceil q\rceil}\left(1-\tfrac{1}{\Delta}\right)\widehat{Q}_{\leq\lceil q\rceil}^{\Lambda}-E$ (3.23)
$\geq\left(1-\tfrac{1}{\Delta}\right)\left(Q_{\leq\lceil q\rceil}^{\Lambda}\left(\mathcal{W}^{\Lambda}+{\lceil q\rceil}-(q+\tfrac{1}{4})\right)+\bar{Q}_{\leq\lceil q\rceil}^{\Lambda}\mathcal{W}^{\Lambda}\left(1-\tfrac{q+\frac{1}{4}}{\lceil q\rceil+1}\right)\right)$
$\geq\tfrac{1}{4\left({\lceil q\rceil}+1\right)}\left(1-\tfrac{1}{\Delta}\right)\mathcal{W}^{\Lambda}.$
It follows from (3.22) that we have a Combes-Thomas bound similar to (3.14) for the modified Green functions: For all $q\in\frac{1}{2}\mathbb{N}^{0}$ we have
$\sup_{z\in\mathds{H}_{q}}\left\lvert\widehat{G}^{\Lambda}_{q,z}({\bf x},{\bf y})\right\rvert\leq C_{q}\mathrm{e}^{-c_{q}\widetilde{d}_{1}^{\Lambda}(\mathbf{x},\mathbf{y})}\quad\text{for all}\quad{\mathbf{x},\mathbf{y}\in\mathcal{P}_{+}(\Lambda).}$ (3.24)
@@
$\pi^{S}_{\nu}=\lceil q\rceil\left(1-\tfrac{1}{\Delta}\right)\widehat{R}^{S}_{q,\nu}\widehat{Q}^{S}_{\leq\lceil q\rceil}\pi^{S}_{\nu}=\lceil q\rceil^{2}\left(1-\tfrac{1}{\Delta}\right)^{2}\widehat{R}^{S}_{q,\nu}\widehat{Q}^{S}_{\leq\lceil q\rceil}\pi^{S}_{\nu}\widehat{R}^{S}_{q,\nu}\widehat{Q}^{S}_{\leq\lceil q\rceil}.$ (3.25)
The proof of Theorem 2.4 will be facilitated by the following a-priori estimate (see, e.g., [12, Lemma 3.4]):
@@
where $\left\lVert\cdot\right\rVert_{2}$ denotes the Hilbert-Schmidt norm, which implies that
@@
In this Section we prove Theorem 2.4. We fix $\Delta>1$, $\lambda>0$, and $s\in(0,\frac{1}{3})$. For $k\in\left\{1,2,\ldots,N\right\}$ we let
@@
[…]
@@
$\sup_{z\in\mathds{H}_{q}}\sup_{\Lambda\subset\mathbb{Z}}\mathbb{E}_{\Lambda}\left\{\left\lvert G^{\Lambda}_{z}({\bf x},{\bf y})\right\rvert^{s}\right\}$ $\leq C_{q}\mathrm{e}^{-c_{q}d_{1}^{\Lambda}(\mathbf{x},\mathbf{y})}+C_{q}\sum_{{\bf u}\in\mathcal{P}_{N,\lceil q\rceil}(\Lambda)}\mathrm{e}^{-c_{q}d_{1}^{\Lambda}(\mathbf{x},\mathbf{u})}\mathrm{e}^{-c_{q}\mathrm{d}_{1}^{\Lambda}(\mathbf{u},\mathbf{y})}$ (4.4)
$\quad\quad+C_{q}\sum_{{\bf u},{\bf v}\in\mathcal{P}_{N,\lceil q\rceil}(\Lambda)}\mathrm{e}^{-c_{q}d_{1}^{\Lambda}(\mathbf{x},\mathbf{u})}\mathrm{e}^{-c_{q}d^{\Lambda}_{H}\left({\bf u},{\bf v}\right)}\mathrm{e}^{-c_{q}d_{1}^{\Lambda}(\mathbf{v},\mathbf{y})}$
@@
where in the last step we used properties of exponential sums (see (B.2)). ∎
@@
We first prove the lemma for $\mathbf{x},\mathbf{y}\in\mathcal{P}_{N,\lceil q\rceil}(D)$. This is done using the reduction to resolvents achieved by using the estimate [4, Eq. (7.44)] and the spectral averaging as in [3, Theorem 4.5]. The final result can be re-formulated in our setting as:
@@
$\mathbb{E}_{D}\left\{\mathcal{Q}_{I}^{D}({\bf x},{\bf y})\right\}\leq C_{r}\sum_{{\bf u}\in\mathcal{P}_{N}(D):x\in{\bf u}}\int_{I}\mathbb{E}_{D}\left\{\left\lvert G_{E}^{D}({\bf u},{\bf y})\right\rvert^{r}\right\}dE\quad\text{for any}\quad x\in{\bf x}.$ (4.8)
Note that that $d^{D}_{H}({\bf u},{\bf y})\geq d^{D}_{H}({x},{\bf y})$ if $x\in\mathbf{u}\subset D$.
Given $\mathbf{x},\mathbf{y}\in\mathcal{P}_{N,\lceil q\rceil}(D)$, we estimate $\mathbb{E}_{D}\left\{\mathcal{Q}_{q}^{D}({\bf x},{\bf y})\right\}$ by (4.8), and estimate the term $\mathbb{E}_{D}\left\{\left\lvert G_{E}^{D}({\bf u},{\bf y})\right\rvert^{s}\right\}$ inside the integral as in (4.4), using (4.5) , getting
$\mathbb{E}_{D}\left\{\mathcal{Q}_{q}^{D}({\bf x},{\bf y})\right\}$ $\leq C_{q}\sum_{{\bf u}\in\mathcal{P}_{N}(D):x\in{\bf u}}\mathrm{e}^{-c_{q}d_{1}^{\mathcal{D}}(\mathbf{u},\mathbf{y})}+C_{q}\sum_{{\bf u}\in\mathcal{P}_{N}(D):x\in{\bf u}}\ \sum_{{\bf v}\in\mathcal{P}_{N,\lceil q\rceil}(D)}\mathrm{e}^{-c_{q}d_{1}^{\mathcal{D}}(\mathbf{u},\mathbf{v})}\mathrm{e}^{-c_{q}\mathrm{d}_{1}^{\mathcal{D}}(\mathbf{v},\mathbf{y})}$ (4.9)
$\quad\quad+C_{q}\sum_{{\bf u}\in\mathcal{P}_{N}(D):x\in{\bf u}}\ \sum_{{\bf v},{\bf w}\in\mathcal{P}_{N,\lceil q\rceil}(D)}\mathrm{e}^{-c_{q}d_{1}^{\mathcal{D}}(\mathbf{u},\mathbf{v})}\mathrm{e}^{-c_{q}d^{D}_{H}\left({\bf v},{\bf w}\right)}\mathrm{e}^{-c_{q}\mathrm{d}_{1}^{\mathcal{D}}({\bf w},\mathbf{y})}.$
To bound the first sum, we note that it follows from (4.7) that
$d_{1}^{\mathcal{D}}(\mathbf{u},\mathbf{y})\geq d_{H}^{D}({{\bf u},{\bf y}})\geq\operatorname{dist}_{D}(x,{\bf y})=d_{H}^{D}({\bf x},{\bf y})\;\text{if}\;x\in\mathbf{u}.$ (4.10)
Hence
$\sum_{{\bf u}\in\mathcal{P}_{N}(D):x\in{\bf u}}\mathrm{e}^{-c_{q}d_{1}^{\mathcal{D}}(\mathbf{u},\mathbf{y})}\leq\mathrm{e}^{-\frac{c_{q}}{2}d_{H}^{D}({\bf x},{\bf y})}\sum_{{\bf u}\in\mathcal{P}_{N}(D)}\mathrm{e}^{-\frac{c_{q}}{2}\mathrm{d}_{1}^{\mathcal{D}}(\mathbf{u},\mathbf{y})}\leq C_{q}\mathrm{e}^{-\frac{c_{q}}{2}d_{H}^{D}({\bf x},{\bf y})},$ (4.11)
@@
$d_{1}^{\mathcal{D}}(\mathbf{u},\mathbf{v})+d_{1}^{\mathcal{D}}(\mathbf{v},\mathbf{y})$ $\geq d_{1}^{\mathcal{D}}(\mathbf{u},\mathbf{y})\geq d_{H}^{D}({\bf x},{\bf y})\quad\text{if}\quad x\in\mathbf{u},$ (4.12)
$d_{1}^{\mathcal{D}}(\mathbf{u},\mathbf{v})+d_{1}^{\mathcal{D}}(\mathbf{v},\mathbf{y})$ $\geq\tfrac{1}{2}\left(d_{1}^{\mathcal{D}}(\mathbf{u},\mathbf{y})+\mathrm{d}_{1}^{\mathcal{D}}(\mathbf{v},\mathbf{y})\right).$
@@
$\sum_{{\bf u}\in\mathcal{P}_{N}(D):x\in{\bf u}}\ \sum_{{\bf v}\in\mathcal{P}_{N,\lceil q\rceil}(D)}\mathrm{e}^{-c_{q}d_{1}^{\mathcal{D}}(\mathbf{u},\mathbf{v})}\mathrm{e}^{-c_{q}d_{1}^{\mathcal{D}}(\mathbf{v},\mathbf{y})}$ (4.13)
$\quad\leq\mathrm{e}^{-\frac{c_{q}}{2}d_{H}^{D}(({\bf x},{\bf y})}\sum_{{\bf u}\in\mathcal{P}_{N}(D)}\sum_{{\bf v}\in\mathcal{P}_{N,\lceil q\rceil}(D)}\mathrm{e}^{-\frac{c_{q}}{4}d_{1}^{\mathcal{D}}(\mathbf{u},\mathbf{y})}e^{-\frac{c_{q}}{2}d_{1}^{\mathcal{D}}(\mathbf{v},\mathbf{y})}\leq C_{q}e^{-\frac{c_{q}}{2}d_{H}^{D}({\bf x},{\bf y})},$